4x+2y-z=7;
8x-2y+z=17

Answer by ankor@dixie-net.com(22740)

4x+2y-z=7;
8x-2y+z=17
--------------adding eliminates y and z
12x + 0 + 0 = 24
x = 24/12
x = 2
:
Replace x with 2 in both equations
4(2) + 2y - z = 7
8 + 2y - z = 7
2y - z = 7 - 8
2y - z = -1
and
8(2) - 2y + z = 17
-2y + z = 17 - 16
-2y + z = 1
multiply by -1
2y - z = -1
:
you can see these two equation are the same, all we can be sure of is x = 2